OpenStudy (anonymous):

4 = 8(1-1.03^n) Solve for n Can you give me steps too?

OpenStudy (anonymous):

is it \[4=8(1-(1.03)^n)\]?

OpenStudy (anonymous):

Yes Satellite73

OpenStudy (anonymous):

I was told the answer is a nice 2 digit number

OpenStudy (anonymous):

ok, to solve this problem, you first need to both sides by 8. After this the equation will look like this. \[.5=(1-1.03^n)\] now, you can subtract the 1 from both sides -.5=-1.03^n Now take the natural log, ln, of both sides. This will give the following. ln(.5)=n*(ln(1.03)) divide ln(.5) by ln(1.03) and there is your answer!

OpenStudy (anonymous):

divide by 8, subtract 1 get \[-\frac{1}{2}=-(1.03)^n\] or \[\frac{1}{2}=(1.03)^n\] then what delta said

OpenStudy (anonymous):

its not a nice number, i got -23.45 double check me though

OpenStudy (anonymous):

although i would recommend knowing that to solve for \[b^x=A\] you can go right to \[x=\frac{\ln(A)}{\ln(b)}\] by the "change of base" formula

OpenStudy (anonymous):

Satellite73 I was told the resulting equation should be an exponential equation in the from \[m ^{x} = n\]

OpenStudy (anonymous):

and indeed it is

OpenStudy (anonymous):

I was also told to use trial and error to find the value of n.

OpenStudy (anonymous):

\[(1.03)^n=.5\]

OpenStudy (anonymous):

really? well try it then but i would solve using logs

OpenStudy (anonymous):

Crap this is really confusing and it is the very last question of my unit on Geometrics